OpenStudy (anonymous):

find the discriminant for x^2+4x+20=0. the roots are... find the discriminant for x^2+4x+20=0. the roots are... @Mathematics

OpenStudy (anonymous):

for ax^2+bx+c=0 the disc is b^2-4ac

OpenStudy (anonymous):

=16-80=-64 roots are complex

OpenStudy (mimi_x3):

If D greater than or equal to 0 the roots are real If D less than 0 the roots are not real If D=0 the roots are equal If D is a perfect square, the roots are rational

OpenStudy (anonymous):

where did you get the a,b, and c from?

OpenStudy (anonymous):

a=coefficient of x^2 b=coeff of x c=constant

OpenStudy (mimi_x3):

x^2 + 4x + 20 = 0 a = 1 b = 4 c =20